find the slope of (2,6) and (10, -8)
find the slope and y intercept of 5x= 7y + 63
find the slope and y intercept of 8 - 1/5y =10x
find the slope of (-4, -7) and (-15, -14)

Answer by checkley77(12569)

find the slope of (2,6) and (10, -8)
slope=(-8-6)/(10-2)=-14/8=-7/4 ans.
--------------------------------------------
find the slope and y intercept of 5x= 7y + 63
7y=5x-63
slope=5 & y intercept=-63
---------------------------------------------------
find the slope and y intercept of 8 - 1/5y =10x
-y/5=10x-8
y/5=-10x+8
y=-50x+40
slope=-50 & y intercept=40
---------------------------------------------------------
find the slope of (-4, -7) and (-15, -14)
slope=(-14+7)/(-15+4)=-7/-11=7/11 ans.